OpenStudy (anonymous):

Malin and Shawn begin with the same number, t. Malin subtracts 6 from t and then divides the result by 6. Shawn adds 7 to t and then divides the result by 7. If Malin's final answer is the same as Shawn's final answer, what was the number t? will give medal.

OpenStudy (welshfella):

t - 6 t + 7 ---- = ---- 6 7 solve for t

OpenStudy (anonymous):

so that is what you do to get the answer?

OpenStudy (welshfella):

yes

OpenStudy (anonymous):

so is the answer 2?

OpenStudy (welshfella):

the first step could be to cross multiply 7t - 42 = 6t + 42

OpenStudy (welshfella):

no

OpenStudy (anonymous):

so what is the next step?

OpenStudy (anonymous):

It's 84.

OpenStudy (anonymous):

Don't really know how to explain.

OpenStudy (lucaz):

84, youre right

OpenStudy (welshfella):

7t - 42 = 6t + 42 add -6t to both sides t - 42 = 42 add 42 to both sides t = 84

OpenStudy (welshfella):

@jf04 plz don't give direct answers

OpenStudy (anonymous):

Sorry, I just didn't know how to explain.

OpenStudy (anonymous):

Normally, I do explain.

OpenStudy (welshfella):

well check out my last post remember with equations what ever you do to one side you have to do to the other - its a bit like a chemical balance

OpenStudy (welshfella):

yea ok